# Less is More: One-shot-subgraph Link Prediction on Large-scale Knowledge Graphs

### v1

@@ 1 Introduction
A knowledge graph (KG) is graph-structural data with relational facts (Battaglia et al., 2018; Ji et al., 2020; Chen et al., 2020), based on which, one can conduct link prediction to deduce new facts from existing ones. The typical problem is to find the answer entity for the specific query, e.g., to find the answer Los Angeles to the query (LeBron, lives_in, ?). With continuous advances in recent years, the link prediction on KG has been widely applied in recommendation systems (Cao et al., 2019; Wang et al., 2019), online question answering (Huang et al., 2019), and drug discovery (Yu et al., 2021). The prediction system learns from the local structure of a KG, where existing methods can be generally summarized as two categories: (1) semantic models that implicitly capture the local evidence through learning the low-dimensional embeddings of entities and relations (Bordes et al., 2013; Dettmers et al., 2017; Zhang et al., 2019; Wang et al., 2017); and (2) structural models that explicitly explore the KG’s structure based on relational paths or graphs with recurrent neural networks (RNNs) or graph neural networks (GNNs) (Das et al., 2017; Schlichtkrull et al., 2018; Sadeghian et al., 2019; Vashishth et al., 2019; Teru et al., 2020; Qu et al., 2021; Zhu et al., 2021; Zhang & Yao, 2022). Although achieving leading performances, these structural models suffer from a severe scalability problem as the entire KG has been potentially or progressively taken for prediction. This inefficient manner hinders their application and optimization on large-scale KGs, e.g., OGB (Hu et al., 2020). Thus, it raises an open question: Is all the information necessary for prediction on knowledge graphs? Intuitively, only partial knowledge stored in the human brain is relevant to a given question, which is extracted by recalling and then utilized in the careful thinking procedure. Similarly, generating candidates and then ranking promising ones are common practices in large-scale recommendation systems with millions even billions of users (Cheng et al., 2016; Covington et al., 2016). These facts motivate us to conduct efficient link prediction with an effective sampling mechanism for KGs. In this work, we propose the novel one-shot-subgraph link prediction on a knowledge graph. This idea paves a new way to alleviate the scalability problem of existing KG methods from a data-centric perspective: decoupling the prediction procedure into two steps with a corresponding sampler and a predictor. Thereby, the prediction of a specific query is conducted by (i) fast sampling of one query-dependent subgraph with the sampler and (ii) slow prediction on the subgraph with predictor. Nevertheless, it is non-trivial to achieve efficient and effective link prediction on large-scale KGs due to the two major challenges. (1) Sampling speed and quality: The fast sampling of the one-shot sampler should be capable of covering the essential evidence and potential answers to support the query. (2) Joint optimization: The sampler and predictor should be optimized jointly to avoid trivial solutions and to guarantee the expressiveness and adaptivity of the overall model to a specific KG. To solve these challenges technically, we first implement the one-shot-subgraph link prediction by the non-parametric and computation-efficient Personalized PageRank (PPR), which is capable of effectively identifying the potential answers without requiring learning. With the efficient subgraph-based prediction, we further propose to search the data-adaptive configurations in both data and model spaces. We show it unnecessary to utilize the whole KG in inference; meanwhile, only a relatively small proportion of information (e.g., 10% of entities) is sufficient. Our main contributions are:
@@ 2 Preliminaries
Notations. A knowledge graph is denoted as $\mathcal{G}\!=\!(\mathcal{V},\mathcal{R},\mathcal{E})$, where the entity set $\mathcal{V}$, the relation set $\mathcal{R}$, and factual edge set =$\mathcal{E}\!=\!\{(x,r,v):x,v\!\in\!\mathcal{V},r\!\in\!\mathcal{R}\}$. Here, a fact is formed as a triplet and denoted as $(x,r,v)$. Besides, a sampled subgraph of $\mathcal{G}$ is denoted as $\mathcal{G}_{s}\!=\!(\mathcal{V}_{s},\mathcal{R}_{s},\mathcal{E}_{s})$, satisfying that $\mathcal{V}_{s}\!\subseteq\!\mathcal{V},\mathcal{R}_{s}\!\subseteq\!\mathcal{R},\mathcal{E}_{s}\!\subseteq\!\mathcal{E}$. The atomic problem of link prediction is denoted as a query $(u,q,?)$, i.e., given the query entity $u$ and query relation $q$, to find the answer entity $v$, making $(u,q,v)$ valid. Semantic models encode entities and relations to low-dimensional entity embeddings $\bm{H}_{\mathcal{V}}\!\in\!\mathbb{R}^{|\mathcal{V}|\!\times\!D_{v}}$ and relation embeddings $\bm{H}_{\mathcal{R}}\!\in\!\mathbb{R}^{|\mathcal{R}|\!\times\!D_{r}}$, where $D_{v},\!D_{r}$ are dimensions. A scoring function $f_{\bm{\theta}}$, e.g., TransE (Bordes et al., 2013) or QuatE (Zhang et al., 2019), is necessary here to quantify the plausibility of a query triplet $(u,q,v)$ with the learned embeddings $(\bm{h}_{u},\bm{h}_{q},\bm{h}_{v})$ as $f_{\bm{\theta}}:(\mathbb{R}^{D_{v}},\mathbb{R}^{D_{r}},\mathbb{R}^{D_{v}})\!\mapsto\!\mathbb{R}$. Efficient semantic models aim to reduce the size of entity embeddings. NodePiece (Galkin et al., 2022) proposes an anchor-based approach that obtains fixed-size embeddings as $\mathcal{G}\!\xmapsto{f_{\bm{\theta}}}\!\bm{\hat{H}}_{\mathcal{V}}\!\in\!\mathbb{R}^{N\!\times\!D_{v}}$ and inference as $(\bm{\hat{H}},\mathcal{G})\xmapsto{f_{\bm{\theta}},(u,q)}\hat{\bm{Y}}$, where $\hat{\bm{Y}}$ are the scores of candidate answers, and $N\!\ll\!|\mathcal{V}|$. Designed to reduce the embedding size, NodePiece cannot reduce the graph size for structural models. Structural models are based on relational paths or graphs for link prediction. Wherein the path-based models, e.g., MINERVA (Das et al., 2017), DRUM (Sadeghian et al., 2019), and RNNLogic (Qu et al., 2021), aim to learn probabilistic and logical rules and well capture the sequential patterns in KGs. The graph-based models such as R-GCN (Schlichtkrull et al., 2018) and CompGCN (Vashishth et al., 2019) propagate the low-level entity embeddings among the neighboring entities to obtain high-level embeddings. Recent methods NBFNet (Zhu et al., 2021) and RED-GNN (Zhang & Yao, 2022) progressively propagate from $u$ to its neighborhood in a breadth-first-searching (BFS) manner. Sampling-based structural models adopt graph sampling approaches to decrease the computation complexity, which can be categorized into two-fold as follows. First, subgraph-wise methods such as GraIL (Teru et al., 2020) and CoMPILE (Mai et al., 2021) extract enclosing subgraphs between query entity $u$ and each candidate answer $v$. Second, layer-wise sampling methods extract a subgraph for message propagation in each layer of a model. Wherein designed for node-level tasks on homogeneous graphs, GraphSAGE (Hamilton et al., 2017) and FastGCN (Chen et al., 2018) randomly sample neighbors around the query entity. While the KG sampling methods, e.g., DPMPN (Xu et al., 2019), AdaProp (Zhang et al., 2023c), and AStarNet (Zhu et al., 2023), extract a learnable subgraph in $\ell$-th layer by the GNN model in $\ell$-th layer, coupling the procedures of sampling and prediction.
@@ 4.1 Realization: Three-step prediction with personalized pagerank
Overview. As illustrated in Fig. 1, the three key steps of our method are (1) generating the sampling distribution $\mathbb{P}_{\mathcal{G}}$ by sampler $g_{\bm{\phi}}$, (2) sampling a subgraph from the distribution as $\mathcal{G}_{s}\!\sim\!\mathbb{P}_{\mathcal{G}}$ with top-$K$ entities and edges, and (3) predicting on the subgraph $\mathcal{G}_{s}$ and acquiring the final prediction $\hat{\bm{Y}}$ by predictor $f_{\bm{\theta}}$. The three-step procedure is summarized in Algorithm 1 and elaborated on as follows. Step-1. Generate sampling distribution. Previous studies show that $v$ is generally near to $u$ (Zhu et al., 2021; Zhang & Yao, 2022), and the relational paths connecting $u$ and $v$ that support the query also lie close to $u$ (Das et al., 2017; Sadeghian et al., 2019). To efficiently capture the local evidence of $u$, we choose the heuristic Personalized PageRank (PPR) (Page et al., 1999; Jeh & Widom, 2003) as the sampling indicator. Note that PPR is not only efficient for its non-parametric nature but also query-dependent and local-structure-preserving for its single-source scoring that starts from $u$. Specifically, PPR starts propagation from $u$ to evaluate the importance of each neighbor of $u$ and generates the PageRank scores as the sampling probability that encodes the local neighborhood of the query entity $u$. Besides, it can also preserve the locality and connectivity of subgraphs by leveraging the information from a large neighborhood. Given a query entity $u$, we obtain the probability $\bm{p}\!\in\!\mathbb{R}^{|\mathcal{V}|}$
@@
by iteratively updating the scores up to $K\!=\!100$ steps to approximate the converged scores efficiently. Here, the initial score $\bm{p}^{(0)}\!=\!\bm{s}\!=\!\mathbbm{1}(u)\!\in\!\{0,1\}^{|\mathcal{V}|}$ indicates the query entity $u$ to be explored. The two-dimensional degree matrix $\bm{D}\!\in\!\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ and adjacency matrix $\bm{A}\!\in\!\{0,1\}^{|\mathcal{V}|\times|\mathcal{V}|}$ together work as the transition matrix, wherein $\bm{A}_{ij}\!=\!1$ means an edge $(i,r,j)\!\in\!\mathcal{E}$ and $\bm{D}_{ij}\!=\!\text{degree}(v_{i})$ if $i\!=\!j$ else $\bm{D}_{ij}\!=\!0$. The damping coefficient $\alpha$ ($=0.85$ by default) controls the differentiation degree. Step-2. Extract a subgraph. Based on the PPR scores $\bm{p}$ (Eqn. 2), the subgraph $\mathcal{G}_{s}\!=\!(\mathcal{V}_{s},\mathcal{E}_{s},\mathcal{R}_{s})$ (where $\mathcal{R}_{s}\!=\!\mathcal{R}$) is extracted with the most important entities and edges. Denoting the sampling ratios of entities and edges as $r^{q}_{\mathcal{V}},r^{q}_{\mathcal{E}}\!\in\!(0,1]$ that depend on the query relation $q$, we sample $|\mathcal{V}_{s}|\!=\!r^{q}_{\mathcal{V}}\!\times\!|\mathcal{V}|$ entities and $|\mathcal{E}_{s}|\!=\!r^{q}_{\mathcal{E}}\!\times\!|\mathcal{E}|$ edges from the full graph $\mathcal{G}$. With the $\texttt{TopK}(D,P,K)$ operation that picks up top-$K$ elements from candidate $D$ w.r.t. probability $P$, the entities $\mathcal{V}_{s}$ and edges $\mathcal{E}_{s}$ are given as
@@
Step-3. Reason on the subgraph. From the model’s perspective, we build the configuration space of the predictor and further utilize the advantages of existing structural models introduced in Sec. 2. Three query-dependent message functions MESS(normal-$\cdot$) are considered, including DRUM, NBFNet, and RED-GNN, which are elaborated in Appendix. B. Note the effective message is propagated from $u$ to the sampled entities $o\in\mathcal{V}_{s}$. Generally, the layer-wise updating of representations is formulated as
@@ 4.2 Optimization: Efficient searching for data-adaptive configurations
Search space. Note that hyperparameters $(r^{q}_{\mathcal{V}},r^{q}_{\mathcal{E}})$ and $L$ play important roles in Algorithm 1. Analytically, a larger subgraph with larger $r^{q}_{\mathcal{V}},r^{q}_{\mathcal{E}}$ does not indicate a better performance, as more irrelevant information is also covered. Besides, a deeper model with a larger $L$ can capture more complex patterns but is more likely to suffer from the over-smoothing problem (Oono & Suzuki, 2019). Overall, the $(r^{q}_{\mathcal{V}},r^{q}_{\mathcal{E}})$ are for sampler’s hyper-parameters $\bm{\phi}_{\text{hyper}}$. In addition to $L$, predictor’s hyper-parameters $\bm{\theta}_{\text{hyper}}$ contain several intra-layer or inter-layer designs, as illustrated in Fig. 2(a). Search problem. Next, we propose the bi-level optimization problem to adaptively search for the optimal configuration $(\bm{\phi}^{*}_{\text{hyper}},\bm{\theta}^{*}_{\text{hyper}})$ of design choices on a specific KG, namely,
@@
where the performance measurement $\mathcal{M}$ can be Mean Reciprocal Ranking (MRR) or Hits@k. Note the non-parametric sampler $g_{\bm{\phi}}$ only contains hyper-parameters $\bm{\phi}_{\text{hyper}}$. As for predictor $f_{\bm{\theta}}$, its $\bm{\theta}\!=\!(\bm{\theta}_{\text{hyper}},\bm{\theta}_{\text{learn}})$ also includes learnable $\bm{\theta}_{\text{learn}}$ that $\bm{\theta}_{\text{learn}}^{*}\!=\!\arg\min_{\bm{\theta}_{\text{learn}}}\mathcal{L}_{cls}(f_{(\bm{\theta}_{\text{hyper}},\bm{\theta}_{\text{learn}})},g_{\bar{\bm{\phi}}_{\text{hyper}}},\mathcal{E}^{\text{train}})$. Search algorithm. Directly searching on both data and model spaces is expensive due to the large space size and data scale. Hence, we split the search into two sub-processes as Fig. 2(b), i.e.,
@@
Specifically, we follow the sequential model-based Bayesian Optimization (BO) (Bergstra et al., 2013; Hutter et al., 2011) to obtain $\bm{\phi}^{*}_{\text{hyper}}$ and $\bm{\theta}^{*}_{\text{hyper}}$. Random forest (RF) (Breiman, 2001) is chosen as the surrogate model because it has a stronger power for approximating the complex and discrete curvature (Grinsztajn et al., 2022), compared with other common surrogates, e.g., Gaussian Process (GP) (Williams & Rasmussen, 1995) or Multi-layer Perceptron (MLP) (Gardner & Dorling, 1998). Acceleration in searching. We adopt a data split trick that balances observations and predictions. It saves time as the training does not necessarily traverse all the training facts. Specifically, partial training facts are sampled as training queries while the others are treated as observations, i.e., we randomly separate the training facts into two parts as $\mathcal{E}^{\text{train}}\!=\!\mathcal{E}^{\text{obs}}\cup\mathcal{E}^{\text{query}}$, where the overall prediction system $f_{\bm{\theta}}\circ g_{\bm{\phi}}$ takes $\mathcal{E}^{\text{obs}}$ as input and then predicts $\mathcal{E}^{\text{query}}$ (see Fig. 2(c)). Here, the split ratio $r^{\text{split}}$ is to balance the sizes of these two parts as $r^{\text{split}}\!=\!\nicefrac{{|\mathcal{E}^{\text{obs}}|}}{{|\mathcal{E}^{\text{query}}|}}$. Thus, the training becomes $\bm{\theta}^{*}\!=\!\arg\min_{\bm{\theta}}\Sigma_{(u,q,v)\in\mathcal{E}^{\text{query}}}\mathcal{L}_{cls}\big{(}f_{\bm{\theta}}(\mathcal{G}_{s}),v\big{)}$ with the split query edges $\mathcal{E}^{\text{query}}$, where $\mathcal{G}_{s}\!=\!g_{\bm{\phi}}(\mathcal{E}^{\text{obs}},u,q)$ with the split observation edges $\mathcal{E}^{\text{obs}}$. More technical details can be found in the Appendix. B.

### v2

# Less is More: One-shot Subgraph Reasoning on Large-scale Knowledge Graphs

@@ 1 Introduction
A knowledge graph (KG) is graph-structural data with relational facts (Battaglia et al., 2018; Ji et al., 2020; Chen et al., 2020), based on which, one can conduct link prediction to deduce new facts from existing ones. The typical problem is to find the answer entity for the specific query, e.g., to find the answer Los Angeles to the query (LeBron, lives_in, ?). With continuous advances in recent years, the link prediction on KG has been widely applied in recommendation systems (Cao et al., 2019; Wang et al., 2019), online question answering (Huang et al., 2019), and drug discovery (Yu et al., 2021).
The prediction system learns from the local structure of a KG, where existing methods can be generally summarized as two categories: (1) semantic models that implicitly capture the local evidence through learning the low-dimensional embeddings of entities and relations (Bordes et al., 2013; Dettmers et al., 2017; Zhang et al., 2019; Wang et al., 2017); and (2) structural models that explicitly explore the KG’s structure based on relational paths or graphs with recurrent neural networks (RNNs) or graph neural networks (GNNs) (Das et al., 2017; Schlichtkrull et al., 2018; Sadeghian et al., 2019; Vashishth et al., 2019; Teru et al., 2020; Qu et al., 2021; Zhu et al., 2021; Zhang & Yao, 2022).
Although achieving leading performances, these structural models suffer from a severe scalability problem as the entire KG has been potentially or progressively taken for prediction. This inefficient manner hinders their application and optimization on large-scale KGs, e.g., OGB (Hu et al., 2020). Thus, it raises an open question: Is all the information necessary for prediction on knowledge graphs? Intuitively, only partial knowledge stored in the human brain is relevant to a given question, which is extracted by recalling and then utilized in the careful thinking procedure. Similarly, generating candidates and then ranking promising ones are common practices in large-scale recommendation systems with millions even billions of users (Cheng et al., 2016; Covington et al., 2016). These facts motivate us to conduct efficient link prediction with an effective sampling mechanism for KGs.
In this work, we propose the novel one-shot-subgraph link prediction on a knowledge graph. This idea paves a new way to alleviate the scalability problem of existing KG methods from a data-centric perspective: decoupling the prediction procedure into two steps with a corresponding sampler and a predictor. Thereby, the prediction of a specific query is conducted by (i) fast sampling of one query-dependent subgraph with the sampler and (ii) slow prediction on the subgraph with predictor.
Nevertheless, it is non-trivial to achieve efficient and effective link prediction on large-scale KGs due to the two major challenges. (1) Sampling speed and quality: The fast sampling of the one-shot sampler should be capable of covering the essential evidence and potential answers to support the query. (2) Joint optimization: The sampler and predictor should be optimized jointly to avoid trivial solutions and to guarantee the expressiveness and adaptivity of the overall model to a specific KG.
To solve these challenges technically, we first implement the one-shot-subgraph link prediction by the non-parametric and computation-efficient Personalized PageRank (PPR), which is capable of effectively identifying the potential answers without requiring learning. With the efficient subgraph-based prediction, we further propose to search the data-adaptive configurations in both data and model spaces. We show it unnecessary to utilize the whole KG in inference; meanwhile, only a relatively small proportion of information (e.g., 10% of entities) is sufficient. Our main contributions are:
@@ 2 Preliminaries
Notations. A knowledge graph is denoted as $\mathcal{G}\!=\!(\mathcal{V},\mathcal{R},\mathcal{E})$, where the entity set $\mathcal{V}$, the relation set $\mathcal{R}$, and factual edge set =$\mathcal{E}\!=\!\{(x,r,v):x,v\!\in\!\mathcal{V},r\!\in\!\mathcal{R}\}$. Here, a fact is formed as a triplet and denoted as $(x,r,v)$. Besides, a sampled subgraph of $\mathcal{G}$ is denoted as $\mathcal{G}_{s}\!=\!(\mathcal{V}_{s},\mathcal{R}_{s},\mathcal{E}_{s})$, satisfying that $\mathcal{V}_{s}\!\subseteq\!\mathcal{V},\mathcal{R}_{s}\!\subseteq\!\mathcal{R},\mathcal{E}_{s}\!\subseteq\!\mathcal{E}$. The atomic problem of link prediction is denoted as a query $(u,q,?)$, i.e., given the query entity $u$ and query relation $q$, to find the answer entity $v$, making $(u,q,v)$ valid.
Semantic models encode entities and relations to low-dimensional entity embeddings $\bm{H}_{\mathcal{V}}\!\in\!\mathbb{R}^{|\mathcal{V}|\!\times\!D_{v}}$ and relation embeddings $\bm{H}_{\mathcal{R}}\!\in\!\mathbb{R}^{|\mathcal{R}|\!\times\!D_{r}}$, where $D_{v},\!D_{r}$ are dimensions. A scoring function $f_{\bm{\theta}}$, e.g., TransE (Bordes et al., 2013) or QuatE (Zhang et al., 2019), is necessary here to quantify the plausibility of a query triplet $(u,q,v)$ with the learned embeddings $(\bm{h}_{u},\bm{h}_{q},\bm{h}_{v})$ as $f_{\bm{\theta}}:(\mathbb{R}^{D_{v}},\mathbb{R}^{D_{r}},\mathbb{R}^{D_{v}})\!\mapsto\!\mathbb{R}$.
Efficient semantic models aim to reduce the size of entity embeddings. NodePiece (Galkin et al., 2022) proposes an anchor-based approach that obtains fixed-size embeddings as $\mathcal{G}\!\xmapsto{f_{\bm{\theta}}}\!\bm{\hat{H}}_{\mathcal{V}}\!\in\!\mathbb{R}^{N\!\times\!D_{v}}$ and inference as $(\bm{\hat{H}},\mathcal{G})\xmapsto{f_{\bm{\theta}},(u,q)}\hat{\bm{Y}}$, where $\hat{\bm{Y}}$ are the scores of candidate answers, and $N\!\ll\!|\mathcal{V}|$. Designed to reduce the embedding size, NodePiece cannot reduce the graph size for structural models.
Structural models are based on relational paths or graphs for link prediction. Wherein the path-based models, e.g., MINERVA (Das et al., 2017), DRUM (Sadeghian et al., 2019), and RNNLogic (Qu et al., 2021), aim to learn probabilistic and logical rules and well capture the sequential patterns in KGs. The graph-based models such as R-GCN (Schlichtkrull et al., 2018) and CompGCN (Vashishth et al., 2019) propagate the low-level entity embeddings among the neighboring entities to obtain high-level embeddings. Recent methods NBFNet (Zhu et al., 2021) and RED-GNN (Zhang & Yao, 2022) progressively propagate from $u$ to its neighborhood in a breadth-first-searching (BFS) manner.
Sampling-based structural models adopt graph sampling approaches to decrease the computation complexity, which can be categorized into two-fold as follows. First, subgraph-wise methods such as GraIL (Teru et al., 2020) and CoMPILE (Mai et al., 2021) extract enclosing subgraphs between query entity $u$ and each candidate answer $v$. Second, layer-wise sampling methods extract a subgraph for message propagation in each layer of a model. Wherein designed for node-level tasks on homogeneous graphs, GraphSAGE (Hamilton et al., 2017) and FastGCN (Chen et al., 2018) randomly sample neighbors around the query entity. While the KG sampling methods, e.g., DPMPN (Xu et al., 2019), AdaProp (Zhang et al., 2023c), and AStarNet (Zhu et al., 2023), extract a learnable subgraph in $\ell$-th layer by the GNN model in $\ell$-th layer, coupling the procedures of sampling and prediction.
@@ 4.1 Realization: Three-step prediction with personalized pagerank
Overview. As illustrated in Fig. 1, the three key steps of our method are (1) generating the sampling distribution $\mathbb{P}_{\mathcal{G}}$ by sampler $g_{\bm{\phi}}$, (2) sampling a subgraph from the distribution as $\mathcal{G}_{s}\!\sim\!\mathbb{P}_{\mathcal{G}}$ with top-$K$ entities and edges, and (3) predicting on the subgraph $\mathcal{G}_{s}$ and acquiring the final prediction $\hat{\bm{Y}}$ by predictor $f_{\bm{\theta}}$. The three-step procedure is summarized in Algorithm 1 and elaborated on as follows.
Step-1. Generate sampling distribution. Previous studies show that $v$ is generally near to $u$ (Zhu et al., 2021; Zhang & Yao, 2022), and the relational paths connecting $u$ and $v$ that support the query also lie close to $u$ (Das et al., 2017; Sadeghian et al., 2019). To efficiently capture the local evidence of $u$, we choose the heuristic Personalized PageRank (PPR) (Page et al., 1999; Jeh & Widom, 2003) as the sampling indicator. Note that PPR is not only efficient for its non-parametric nature but also query-dependent and local-structure-preserving for its single-source scoring that starts from $u$.
Specifically, PPR starts propagation from $u$ to evaluate the importance of each neighbor of $u$ and generates the PageRank scores as the sampling probability that encodes the local neighborhood of the query entity $u$. Besides, it can also preserve the locality and connectivity of subgraphs by leveraging the information from a large neighborhood. Given a query entity $u$, we obtain the probability $\bm{p}\!\in\!\mathbb{R}^{|\mathcal{V}|}$
@@
by iteratively updating the scores up to $K\!=\!100$ steps to approximate the converged scores efficiently. Here, the initial score $\bm{p}^{(0)}\!=\!\bm{s}\!=\!\mathbbm{1}(u)\!\in\!\{0,1\}^{|\mathcal{V}|}$ indicates the query entity $u$ to be explored. The two-dimensional degree matrix $\bm{D}\!\in\!\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ and adjacency matrix $\bm{A}\!\in\!\{0,1\}^{|\mathcal{V}|\times|\mathcal{V}|}$ together work as the transition matrix, wherein $\bm{A}_{ij}\!=\!1$ means an edge $(i,r,j)\!\in\!\mathcal{E}$ and $\bm{D}_{ij}\!=\!\text{degree}(v_{i})$ if $i\!=\!j$ else $\bm{D}_{ij}\!=\!0$. The damping coefficient $\alpha$ ($=0.85$ by default) controls the differentiation degree.
Step-2. Extract a subgraph. Based on the PPR scores $\bm{p}$ (Eqn. 2), the subgraph $\mathcal{G}_{s}\!=\!(\mathcal{V}_{s},\mathcal{E}_{s},\mathcal{R}_{s})$ (where $\mathcal{R}_{s}\!=\!\mathcal{R}$) is extracted with the most important entities and edges. Denoting the sampling ratios of entities and edges as $r^{q}_{\mathcal{V}},r^{q}_{\mathcal{E}}\!\in\!(0,1]$ that depend on the query relation $q$, we sample $|\mathcal{V}_{s}|\!=\!r^{q}_{\mathcal{V}}\!\times\!|\mathcal{V}|$ entities and $|\mathcal{E}_{s}|\!=\!r^{q}_{\mathcal{E}}\!\times\!|\mathcal{E}|$ edges from the full graph $\mathcal{G}$. With the $\texttt{TopK}(D,P,K)$ operation that picks up top-$K$ elements from candidate $D$ w.r.t. probability $P$, the entities $\mathcal{V}_{s}$ and edges $\mathcal{E}_{s}$ are given as
@@
Step-3. Reason on the subgraph. From the model’s perspective, we build the configuration space of the predictor and further utilize the advantages of existing structural models introduced in Sec. 2. Three query-dependent message functions MESS($\cdot$) are considered, including DRUM, NBFNet, and RED-GNN, which are elaborated in Appendix. B. Note the effective message is propagated from $u$ to the sampled entities $o\in\mathcal{V}_{s}$. Generally, the layer-wise updating of representations is formulated as
@@ 4.2 Optimization: Efficient searching for data-adaptive configurations
Search space. Note that hyperparameters $(r^{q}_{\mathcal{V}},r^{q}_{\mathcal{E}})$ and $L$ play important roles in Algorithm 1. Analytically, a larger subgraph with larger $r^{q}_{\mathcal{V}},r^{q}_{\mathcal{E}}$ does not indicate a better performance, as more irrelevant information is also covered. Besides, a deeper model with a larger $L$ can capture more complex patterns but is more likely to suffer from the over-smoothing problem (Oono & Suzuki, 2019). Overall, the $(r^{q}_{\mathcal{V}},r^{q}_{\mathcal{E}})$ are for sampler’s hyper-parameters $\bm{\phi}_{\text{hyper}}$. In addition to $L$, predictor’s hyper-parameters $\bm{\theta}_{\text{hyper}}$ contain several intra-layer or inter-layer designs, as illustrated in Fig. 2(a).
Search problem. Next, we propose the bi-level optimization problem to adaptively search for the optimal configuration $(\bm{\phi}^{*}_{\text{hyper}},\bm{\theta}^{*}_{\text{hyper}})$ of design choices on a specific KG, namely,
@@
where the performance measurement $\mathcal{M}$ can be Mean Reciprocal Ranking (MRR) or Hits@k. Note the non-parametric sampler $g_{\bm{\phi}}$ only contains hyper-parameters $\bm{\phi}_{\text{hyper}}$. As for predictor $f_{\bm{\theta}}$, its $\bm{\theta}\!=\!(\bm{\theta}_{\text{hyper}},\bm{\theta}_{\text{learn}})$ also includes learnable $\bm{\theta}_{\text{learn}}$ that $\bm{\theta}_{\text{learn}}^{*}\!=\!\arg\min_{\bm{\theta}_{\text{learn}}}\mathcal{L}_{cls}(f_{(\bm{\theta}_{\text{hyper}},\bm{\theta}_{\text{learn}})},g_{\bar{\bm{\phi}}_{\text{hyper}}},\mathcal{E}^{\text{train}})$.
Search algorithm. Directly searching on both data and model spaces is expensive due to the large space size and data scale. Hence, we split the search into two sub-processes as Fig. 2(b), i.e.,
@@
Specifically, we follow the sequential model-based Bayesian Optimization (BO) (Bergstra et al., 2013; Hutter et al., 2011) to obtain $\bm{\phi}^{*}_{\text{hyper}}$ and $\bm{\theta}^{*}_{\text{hyper}}$. Random forest (RF) (Breiman, 2001) is chosen as the surrogate model because it has a stronger power for approximating the complex and discrete curvature (Grinsztajn et al., 2022), compared with other common surrogates, e.g., Gaussian Process (GP) (Williams & Rasmussen, 1995) or Multi-layer Perceptron (MLP) (Gardner & Dorling, 1998).
Acceleration in searching. We adopt a data split trick that balances observations and predictions. It saves time as the training does not necessarily traverse all the training facts. Specifically, partial training facts are sampled as training queries while the others are treated as observations, i.e., we randomly separate the training facts into two parts as $\mathcal{E}^{\text{train}}\!=\!\mathcal{E}^{\text{obs}}\cup\mathcal{E}^{\text{query}}$, where the overall prediction system $f_{\bm{\theta}}\circ g_{\bm{\phi}}$ takes $\mathcal{E}^{\text{obs}}$ as input and then predicts $\mathcal{E}^{\text{query}}$ (see Fig. 2(c)). Here, the split ratio $r^{\text{split}}$ is to balance the sizes of these two parts as $r^{\text{split}}\!=\!\nicefrac{{|\mathcal{E}^{\text{obs}}|}}{{|\mathcal{E}^{\text{query}}|}}$. Thus, the training becomes $\bm{\theta}^{*}\!=\!\arg\min_{\bm{\theta}}\Sigma_{(u,q,v)\in\mathcal{E}^{\text{query}}}\mathcal{L}_{cls}\big{(}f_{\bm{\theta}}(\mathcal{G}_{s}),v\big{)}$ with the split query edges $\mathcal{E}^{\text{query}}$, where $\mathcal{G}_{s}\!=\!g_{\bm{\phi}}(\mathcal{E}^{\text{obs}},u,q)$ with the split observation edges $\mathcal{E}^{\text{obs}}$. More technical details can be found in the Appendix. B.